A baseball is shut with an initial speed of 12.0m/s.

(a) how far does it go in the horizontal direction if it is shut at angle (I) 15° (ii) 50° (III) 45 and (iv) 60° above the horizontal


1
Expert's answer
2022-02-04T07:42:08-0500

The range is


"R=\dfrac{v^2\sin(2\theta)}{g}."

(I) 15°: 7.3 m.

(ii) 50°: 14.5 m.

(III) 45°: 14.7 m.

(iv) 60°: 12.7 m.